1. 2C4H10 + 13O2  →  10H2O + 8CO2 what is the molar mass of o2? 
  2. 2C4H10 + 13O2  →  10H2O + 8CO2 What is the molar mass of C4H10 ?
  3. 2C4H10 + 13O2  →  10H2O + 8CO2.   If you are given 150 grams of C4H10 ,How many moles would that be? 
  4. 2C4H10 + 13O2  →  10H2O + 8CO2    If you are given 150 grams of C4H10 ,How many moles of O2 would you need to completely react with the C4H10 ? 
  5. 2C4H10 + 13O2  →  10H2O + 8CO2    If you are given 150 grams of C4H10 ,How many grams of O2 would you need to completely react with the C4H10 ? 
1
Expert's answer
2021-02-18T07:46:23-0500

1. Molar mass of O2= 16x2= 32g/mol

2. Molar mass of C4H10= (12x4) + (1x10)=58g/mol

3. Mole= mass/molar mass

= 150/58= 2.6mol


4.2moles of C4H10 reacts witg 13moles of O2

2.6moles will react with 13/2 x 2.6= 16.9m

5.mass of oxygen = mole x molar mass= 16.9 x 32= 540.8g